OpenStudy (anonymous):

Perform indicated operations and simplify.(3yz)/(4x)+(4xz)/(3y)

OpenStudy (anonymous):

Using math identities will make it easy

OpenStudy (anonymous):

(3yz*3y+4xz*4x)/12xy=(9y^2z+16x^2z)/12xy=z(9y^2+16x^2)/12xy